Consider the following information to answer questions Government spending = R550 Exports = R330 Autonomous consumption = R280 Autonomous imports = R170 Investment expenditure = R120 Marginal propensity to consume =0,75 Full employment level of income =R5700

What is the marginal propensity to save?
a. 1
b. 1.60
c. 0.25
d. 0.40
1
Expert's answer
2020-04-22T11:40:52-0400

"MPC+MPS=1."

So, "MPS=1-MPC=1-0.75=0.25."

So, the answer is c)0.25.